OpenStudy (darkbluechocobo):

Calculate the velocity: displacement = 3500km, time = 3hours

geerky42 (geerky42):

You know the formula? \(v = \dfrac{d}{t}\)

geerky42 (geerky42):

do you need unit of v be m/s?

OpenStudy (darkbluechocobo):

hai o.o m&ms ?

OpenStudy (darkbluechocobo):

how would you get 3 hours to seconds

OpenStudy (darkbluechocobo):

324.07 would this be what we are looking for?

OpenStudy (darkbluechocobo):

@ganeshie8

ganeshie8 (ganeshie8):

you had travelled 3500km in 3hours

ganeshie8 (ganeshie8):

the velocity would be \[\large \dfrac{3500}{3} km/hour\]

OpenStudy (darkbluechocobo):

1166.66 km/h

OpenStudy (darkbluechocobo):

oh wait i need this in m/s

OpenStudy (darkbluechocobo):

the answer has to be in m/s

ganeshie8 (ganeshie8):

how many meters are there in a km ?

ganeshie8 (ganeshie8):

3500km = 3500*1000 meters right ?

ganeshie8 (ganeshie8):

3 hours = 3*60*60 seconds

OpenStudy (darkbluechocobo):

well there is 1000 meters in a km

OpenStudy (darkbluechocobo):

and 3 hours is 60 60 60

ganeshie8 (ganeshie8):

velcity = \[\large \dfrac{3500\times 1000}{3\times 60\times 60} ~m/s\]

ganeshie8 (ganeshie8):

simplify

OpenStudy (darkbluechocobo):

324.07 m/s

OpenStudy (darkbluechocobo):

@ganeshie8 I need one more thing :c

ganeshie8 (ganeshie8):

yeah

OpenStudy (darkbluechocobo):